# A Single-Loop Gradient Algorithm for Pessimistic Bilevel Optimization via Smooth Approximation

### v1

@@ 5.3 Hyper-representation
Hyper-representationgrazzi2020iteration ; franceschi2018bilevel aim to learn an effective representation of the input data for lower-level classifiers, where PBO model was used to handle the potential multiplicity of optimal solutions in the lower-level problem and robust learn the representation guanadaprox . In this experiment, we further explore the potential of the PBO model and compare it with optimistic models.
@@ 5.3.1 Linear hyper-representation on synthetic data
$\min_{H\in\mathbb{R}^{n\times p}}\max_{w}\frac{1}{m_{1}}\|X_{val}^{T}Hw-\mathbf{y}_{val}\|^{2},\quad\text{s.t.}\;w\in\underset{w^{\prime}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{m_{2}}\|X_{train}^{T}Hw^{\prime}-\mathbf{y}_{train}\|^{2},$ (18)
where $\mathbf{X}_{val}\in\mathbb{R}^{n\times m_{1}}$ and $\mathbf{X}_{train}\in\mathbb{R}^{n\times m_{2}}$ are the validation and training feature matrices, and $\mathbf{y}_{val}\in\mathbb{R}^{m_{1}},\mathbf{y}_{train}\in\mathbb{R}^{m_{2}}$ are the corresponding response vectors. The synthetic data is generated as in grazzi2020iteration , with feature matrices $\mathbf{X}_{val},\mathbf{X}_{train},\mathbf{X}_{test}$ and ground-truth matrices $H_{real}$ and vectors $w_{real}$ sampled randomly. The response vectors are formed using the linear model $\mathbf{y}_{(\cdot)}=\mathbf{X}_{(\cdot)}^{\top}H_{real}w_{real}$, with Gaussian noise $\epsilon\sim\mathcal{N}(0,a^{2})$ added to both $\mathbf{X}_{(\cdot)}$ and $\mathbf{y}_{(\cdot)}$ for train and valid data to simulate noise.
To evaluate solver efficiency and formulation effectiveness, we conduct a two-part comparison. First, we compare the SiPBA algorithm with PBO algorithms AdaProx-PD and AdaProx-SG. Second, we assess the impact of the bilevel formulation by comparing the pessimistic model (18) (solved by SiPBA) with its optimistic variant (solved by AID-FP , AID-CG grazzi2020iteration and PZOBO sow2022convergence ), which replaces $\underset{w}{\max}$ with $\underset{w}{\min}$ in the upper level. To assess the robustness of each method under varying levels of noise, we conduct experiments with moderate ($a=0.1$) and severe ($a=1$) perturbations. The performance is measured by test loss, averaged over 10 random seeds. Results in Figure 2 demonstrate the stability and efficiency of SiPBA.
@@ 5.3.2 Deep hyper-representation on MNIST and FashionMNIST
$\min_{\theta\in\Theta}\max_{w}\frac{1}{m_{1}}\|f(X_{val},\theta)w-\mathbf{y}_{val}\|^{2},\;\text{s.t.}\;w\in\underset{w^{\prime}\in\mathcal{W}}{\arg\min}\frac{1}{m_{2}}\|f(X_{train},\theta)w^{\prime}-\mathbf{y}_{train}\|^{2},$ (19)
@@
We adopt the LeNet-5 architecture sow2022convergence as the feature extractor $f(\cdot,\theta)$ and evaluate performance on the MNIST and FashionMNIST datasets. Each dataset is randomly split into 50,000 training samples, 10,000 validation samples, and 10,000 test samples, with performance evaluated by test accuracy. We compare the pessimistic formulation (19), trained with SiPBA, to its optimistic variant (replacing $\underset{w}{\max}$ with $\underset{w}{\min}$ in the upper level) trained with AID-FP, AID-CG grazzi2020iteration , and PZOBO sow2022convergence . Mean results over ten runs are shown in Figure 3, which shows that SiPBA achieves the highest test accuracy.
@@ Acknowledgements
Authors listed in alphabetical order. This work was supported by National Key R&D Program of China (2023YFA1011400), National Natural Science Foundation of China (12222106, 12326605), Guangdong Basic and Applied Basic Research Foundation (No. 2022B1515020082), the Longhua District Science and Innovation Commission Project Grants of Shenzhen (Grant No.20250113G43468522) and Natural Science Foundation of Shenzhen (20250530150024003).
@@ B.1 Equivalent minimax reformulation of $\phi(x)$
The $\mu$-strongly convexity of $\psi_{\rho,\sigma}(x,y,z)$ with respect to $y$ ensures the uniqueness of the minimizer of $\min_{y\in Y}-\psi_{\rho,\sigma}(x,y,z)$ (equivalently, the maximizer of $\max_{y\in Y}\psi_{\rho,\sigma}(x,y,z)$). Let us denote this unique maximizer as $\widehat{y}^{*}(x,z)$. Furthermore, it can be shown that $\psi_{\rho,\sigma}(x,\cdot,z)$ satisfies the inf-compactness condition as stated in [14, Theorem 4.13] on any point $(\bar{x},\bar{z})\in X\times Y$. Specifically, for any $(\bar{x},\bar{z})\in X\times Y$, there exists a constant $c\in\mathbb{R}$, a compact set $B\subset\mathbb{R}^{m}$, and a neighborhood $W$ of $(\bar{x},\bar{z})$ such that the level set $\{y\in Y~|~-\psi_{\rho,\sigma}(x,y,z)\leq c\}$ is nonempty and contained in $B$ for all $(x,z)\in W$. Given that $\psi_{\rho,\sigma}(x,y,z)$ is continuously differentiable, $\widehat{y}^{*}(x,z)$ is unique, and the inf-compactness condition holds, we can apply [14, Theorem 4.13, Remark 4.14]. This theorem implies that $h(x,z)$ is differentiable on $X\times Y$, and its gradient is given by:
@@
$F(\bar{x},y_{k}^{*}(\bar{x}))\geq\phi(\bar{x})+\delta.$
Taking $k\rightarrow\infty$ in the above inequality, since $F(x,y)$ is continuous on $X\times Y$, we have
@@ B.5 Lower semi-continuity of $\phi(x)$
First, let $\tilde{y}_{k}\in\mathcal{S}(x_{k})$ be a sequence converging to $\tilde{y}$, then we assert that $\tilde{y}\in\mathcal{S}(\bar{x})$. Assume by contradiction that $\tilde{y}\notin\mathcal{S}(\bar{x})$. Then there exist a $\delta>0$ such that
$f(\bar{x},y^{*}(\bar{x}))<f(\bar{x},\tilde{y})-\delta.$
Then by the continuity of $f$ and convergence of $x_{k}$, for sufficiently large $k$,
$f(x_{k},\tilde{y}_{k})>f(\bar{x},\tilde{y})-\frac{\delta}{2}>f(\bar{x},y^{*}(\bar{x}))+\frac{\delta}{2}>f(x_{k},y^{*}(\bar{x})).$
which contradicts with $\tilde{y}_{k}\in\mathcal{S}(x_{k})$. Thus, we have
$\underset{x\to\bar{x}}{\limsup}\;\mathcal{S}(x)\in\mathcal{S}(\bar{x}),$
where $\limsup\limits_{x\to\bar{x}}\mathcal{S}(x):=\{y\;|\;\exists x_{k}\to\bar{x},\;y_{k}\in\mathcal{S}(x_{k}),\;\text{s.t.}\;y_{k}\to y\}$. As a result, recalling the definition of the value function $\phi(x)=\max_{y\in\mathcal{S}(x)}F(x,y)$, we have
$\phi(\bar{x})\geq\limsup_{x\to\bar{x}}\phi(x),$
which implies that $\phi(\bar{x})$ is upper semi-continuous. Combining the assumption that $\phi(\bar{x})$ is lower semi-continuous, we have $\phi(\bar{x})$ is continuous. As a result, there exist a sequence $\tilde{y}_{k}:=y^{*}(x_{k})\in\mathcal{S}(x_{k})$ such that $\lim_{k\to\infty}\tilde{y}_{k}=y^{*}(\bar{x}).$ Since $(y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))$ is the saddle point, we have
$\psi_{k}(x_{k},y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))\geq\psi_{k}(x_{k},\tilde{y}_{k},z_{k}^{*}(x_{k})).$
Recall the definition of $\psi_{k}$ and rearrange the items,
$\lim_{k\to\infty}f(x_{k},y_{k}^{*}(x_{k}))-f(x_{k},\tilde{y}_{k})$
$\leq$ $\lim_{k\to\infty}\frac{1}{\rho_{k}}\left(F(x_{k},y_{k}^{*}(x_{k}))-F(x_{k},\tilde{y}_{k})\right)-\frac{\sigma_{k}}{\rho_{k}}\left(\langle y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k})\rangle-\langle\tilde{y}_{k},z_{k}^{*}(x_{k})\rangle\right)$
$=$ $0,$
where the last equation follows from the boundedness of $(y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))$ established in Lemma B.2. Thus, we have any accumulation point of $y_{k}^{*}(x_{k})$ belongs to $\mathcal{S}(\bar{x})$. Furthermore, by rearranging the items above again,
$\lim_{k\to\infty}F(x_{k},y_{k}^{*}(x_{k}))-F(x_{k},\tilde{y}_{k})$
$\geq$ $\lim_{k\to\infty}\rho_{k}\left(f(x_{k},y_{k}^{*}(x_{k}))-f(x_{k},\tilde{y}_{k})\right)+\sigma_{k}\left(\langle y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k})\rangle-\langle\tilde{y}_{k},z_{k}^{*}(x_{k})\rangle\right)$
$\geq$ $0,$
which implies that $\lim_{k\to\infty}F(x_{k},y_{k}^{*}(x_{k}))\geq\lim_{k\to\infty}F(x_{k},\tilde{y}_{k})=F(\bar{x},y^{*}(\bar{x}))$. Combining this with the fact that any accumulation point of $y_{k}^{*}(x_{k})$ belongs to $\mathcal{S}(\bar{x})$, we can conclude that $\lim_{k\to\infty}y_{k}^{*}(x_{k})=y^{*}(\bar{x})$. Furthermore, the saddle point property implies that
$\psi_{k}(x_{k},y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))\leq\psi_{k}(x_{k},y_{k}^{*}(x_{k}),\tilde{y}_{k}).$
By expanding $\psi_{k}$ and rearranging the items, we have
$\lim_{k\to\infty}\|z_{k}^{*}(x_{k})-y_{k}^{*}(x_{k})\|^{2}$
$\leq$ $\lim_{k\to\infty}\frac{\rho_{k}}{\sigma_{k}}(f(x_{k},\tilde{y}_{k})-f(x_{k},z_{k}^{*}(x_{k})))+\|\tilde{y}_{k}-y_{k}^{*}(x_{k})\|^{2}$
$\leq$ $0.$
Combining this with $\lim_{k\to\infty}y_{k}^{*}(x_{k})=y^{*}(\bar{x})$, the conclusion follows. $\square$

### v4

@@ 5.3 Hyper-representation
Hyper-representationGrazzi et al. (2020); Franceschi et al. (2018) aim to learn an effective representation of the input data for lower-level classifiers, where PBO model was used to handle the potential multiplicity of optimal solutions in the lower-level problem and robust learn the representation Guan et al. (2025). In this experiment, we further explore the potential of the PBO model and compare it with optimistic models.
@@ 5.3.1 Linear hyper-representation on synthetic data
$\min_{H\in\mathbb{R}^{n\times p}}\max_{w}\frac{1}{m_{1}}\|\mathbf{X}_{val}^{T}Hw-\mathbf{y}_{val}\|^{2},\quad\text{s.t.}\;w\in\underset{w^{\prime}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{m_{2}}\|\mathbf{X}_{train}^{T}Hw^{\prime}-\mathbf{y}_{train}\|^{2},$ (18)
where $\mathbf{X}_{val}\in\mathbb{R}^{n\times m_{1}}$ and $\mathbf{X}_{train}\in\mathbb{R}^{n\times m_{2}}$ are the validation and training feature matrices, and $\mathbf{y}_{val}\in\mathbb{R}^{m_{1}},\mathbf{y}_{train}\in\mathbb{R}^{m_{2}}$ are the corresponding response vectors. The synthetic data is generated as in Grazzi et al. (2020), with feature matrices $\mathbf{X}_{val},\mathbf{X}_{train},\mathbf{X}_{test}$ and ground-truth matrices $H_{real}$ and vectors $w_{real}$ sampled randomly. The response vectors are formed using the linear model $\mathbf{y}_{(\cdot)}=\mathbf{X}_{(\cdot)}^{\top}H_{real}w_{real}$, with Gaussian noise $\epsilon\sim\mathcal{N}(0,a^{2})$ added to both $\mathbf{X}_{(\cdot)}$ and $\mathbf{y}_{(\cdot)}$ for train and valid data to simulate noise.
To evaluate solver efficiency and formulation effectiveness, we conduct a two-part comparison. First, we compare the SiPBA algorithm with PBO algorithms AdaProx-PD and AdaProx-SG. Second, we assess the impact of the bilevel formulation by comparing the pessimistic model (18) (solved by SiPBA) with its optimistic variant (solved by AID-FP , AID-CG Grazzi et al. (2020) and PZOBO Sow et al. (2022)), which replaces $\underset{w}{\max}$ with $\underset{w}{\min}$ in the upper level. To assess the robustness of each method under varying levels of noise, we conduct experiments with moderate ($a=0.1$) and severe ($a=1$) perturbations. The performance is measured by test loss, averaged over 10 random seeds. Results in Figure 2 demonstrate the stability and efficiency of SiPBA.
@@ 5.3.2 Deep hyper-representation on MNIST and FashionMNIST
$\min_{\theta\in\Theta}\max_{w}\frac{1}{m_{1}}\|f(\mathbf{X}_{val},\theta)w-\mathbf{y}_{val}\|^{2},\;\text{s.t.}\;w\in\underset{w^{\prime}\in\mathcal{W}}{\arg\min}\frac{1}{m_{2}}\|f(\mathbf{X}_{train},\theta)w^{\prime}-\mathbf{y}_{train}\|^{2},$ (19)
@@
We adopt the LeNet-5 architecture Sow et al. (2022) as the feature extractor $f(\cdot,\theta)$ and evaluate performance on the MNIST and FashionMNIST datasets. Each dataset is randomly split into 50,000 training samples, 10,000 validation samples, and 10,000 test samples, with performance evaluated by test accuracy. We compare the pessimistic formulation (19), trained with SiPBA, to its optimistic variant (replacing $\underset{w}{\max}$ with $\underset{w}{\min}$ in the upper level) trained with AID-FP, AID-CG Grazzi et al. (2020), and PZOBO Sow et al. (2022). Mean results over ten runs are shown in Figure 3, which shows that SiPBA achieves the highest test accuracy.
@@ Acknowledgements
Authors listed in alphabetical order. This work was supported by National Key R&D Program of China (2023YFA1011400), National Natural Science Foundation of China (12222106, 12326605), Guangdong Basic and Applied Basic Research Foundation (No. 2022B1515020082), the Longhua District Science and Innovation Commission Project Grants of Shenzhen (No. 20250113G43468522) and Natural Science Foundation of Shenzhen (No. 20250530150024003).
@@ B.1 Equivalent minimax reformulation of $\phi(x)$
The $\mu$-strongly convexity of $\psi_{\rho,\sigma}(x,y,z)$ with respect to $y$ ensures the uniqueness of the minimizer of $\min_{y\in Y}-\psi_{\rho,\sigma}(x,y,z)$ (equivalently, the maximizer of $\max_{y\in Y}\psi_{\rho,\sigma}(x,y,z)$). Let us denote this unique maximizer as $\widehat{y}^{*}(x,z)$. Furthermore, it can be shown that $\psi_{\rho,\sigma}(x,\cdot,z)$ satisfies the inf-compactness condition as stated in [14, Theorem 4.13] on any point $(\bar{x},\bar{z})\in X\times Y$. Specifically, for any $(\bar{x},\bar{z})\in X\times Y$, there exists a constant $c\in\mathbb{R}$, a compact set $B\subset\mathbb{R}^{m}$, and a neighborhood $W$ of $(\bar{x},\bar{z})$ such that the level set $\{y\in Y~|~-\psi_{\rho,\sigma}(x,y,z)\leq c\}$ is nonempty and contained in $B$ for all $(x,z)\in W$.
Given that $\psi_{\rho,\sigma}(x,y,z)$ is continuously differentiable, $\widehat{y}^{*}(x,z)$ is unique, and the inf-compactness condition holds, we can apply [14, Theorem 4.13, Remark 4.14]. This theorem implies that $h(x,z)$ is differentiable on $X\times Y$, and its gradient is given by:
@@
$F(\bar{x},y_{k}^{*}(\bar{x}))-\frac{\sigma_{k}}{2}\|y_{k}^{*}(\bar{x})\|^{2}\geq\phi(\bar{x})+\delta.$
Taking $k\rightarrow\infty$ in the above inequality, since $F(x,y)$ is continuous on $X\times Y$, $\{y_{k}^{*}(\bar{x})\}$ is bounded and $\sigma_{k}\rightarrow 0$, we have
@@ B.5 Lower semi-continuity of $\phi(x)$
Since $\{x_{k}\}$ is bounded, Lemmas B.2 and B.4 imply that the sequences $\{y^{*}(x_{k})\}$ and $\{(y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))\}$ are also bounded.
First, we show that $\mathcal{S}(x)$ is outer semi-continuous on $X$.
[…]
$).f(x_{j},y_{j})\leq f(x_{j},y^{*}(\tilde{x}$)).
Taking the limit as $j\to\infty$ and using the continuity of $f$ on $X\times Y$, we obtain
$f(\tilde{x},\tilde{y})\leq f(\tilde{x},y^{*}(\tilde{x}))=\min_{y\in Y}f(\tilde{x},y).$
[…]
Next, we establish that
$\lim_{k\to\infty}y^{*}(x_{k})=y^{*}(\bar{x}).$
Let $\{y^{*}(x_{j})\}$ be any subsequence of $\{y^{*}(x_{k})\}$ such that $y_{j}^{*}\rightarrow\bar{y}$ as $j\rightarrow\infty$. By assumption, $x_{j}\rightarrow\bar{x}$. From the outer semi-continuity of $\mathcal{S}(x)$ established above, it follows that $\bar{y}\in\mathcal{S}(\bar{x})$. Moreover, by definition $\phi(x_{j})=F(x_{j},y^{*}(x_{j}))$ and $\phi(\bar{x})=F(\bar{x},y^{*}(\bar{x}))$. Using the continuity of $F$ on $X\times Y$ and the lower semi-continuity of $\phi$, we obtain
$F(\bar{x},\bar{y})=\lim_{j\to\infty}F(x_{j},y^{*}(x_{j}))=\lim_{j\to\infty}\phi(x_{j})\geq\phi(\bar{x})=F(\bar{x},y^{*}(\bar{x})).$
Since $\bar{y}\in\mathcal{S}(\bar{x})$, the above inequality implies that $\bar{y}\in\arg\max_{y\in\mathcal{S}(\bar{x})}F(\bar{x},y)$. Because $\mathcal{S}(\bar{x})$ is convex, and $F(\bar{x},y)$ is strongly concave in $y$, this maximizer is unique, so $\bar{y}=y^{*}(\bar{x})$. Hence any accumulation point of sequence $\{y^{*}(x_{k})\}$ equals $y^{*}(\bar{x})$. Since $\{y^{*}(x_{k})\}$ is bounded, we conclude that $\lim_{k\to\infty}y^{*}(x_{k})=y^{*}(\bar{x})$.
Third, we show that
$\lim_{k\to\infty}y_{k}^{*}(x_{k})=y^{*}(\bar{x}).$
Let $\{y_{j}^{*}(x_{j})\}$ be any subsequence of $\{y_{k}^{*}(x_{k})\}$ such that $y_{j}^{*}(x_{j})\rightarrow\bar{y}$ as $j\rightarrow\infty$. By Lemma B.3, we have $\bar{y}\in\mathcal{S}(\bar{x})$. Since $y_{j}^{*}(x_{j})$ maximizes $\psi_{j}(x_{j},y,z_{j}^{*}(x_{j}))$ over $y\in Y$, it follows that
$\psi_{j}(x_{j},y_{j}^{*}(x_{j}),z_{j}^{*}(x_{j}))\geq\psi_{j}(x_{j},y^{*}(x_{j}),z_{j}^{*}(x_{j})).$
Expanding both sides gives
$F(x_{j},y_{j}^{*}(x_{j}))-F(x_{j},y^{*}(x_{j}))\geq$ $\rho_{j}\left(f(x_{j},y_{j}^{*}(x_{j}))-f(x_{j},y^{*}(x_{j}))\right)$
$+\sigma_{j}\left(\langle y_{j}^{*}(x_{j}),z_{j}^{*}(x_{j})\rangle-\langle y^{*}(x_{j}),z_{j}^{*}(x_{j})\rangle\right).$
Since $y^{*}(x_{j})\in\mathcal{S}(x_{j})$, we have $f(x_{j},y_{j}^{*}(x_{j}))-f(x_{j},y^{*}(x_{j}))\geq 0$ and thus
$F(x_{j},y_{j}^{*}(x_{j}))-F(x_{j},y^{*}(x_{j}))\geq\sigma_{j}\left(\langle y_{j}^{*}(x_{j}),z_{j}^{*}(x_{j})\rangle-\langle y^{*}(x_{j}),z_{j}^{*}(x_{j})\rangle\right)$
Taking the limit as $j\rightarrow\infty$ and using the continuity of $F$, together with $\lim_{k\to\infty}y^{*}(x_{k})=y^{*}(\bar{x})$, yields
$F(\bar{x},\bar{y})-F(\bar{x},y^{*}(\bar{x}))\geq 0.$
Since $\bar{y}\in\mathcal{S}(\bar{x})$, this implies that $\bar{y}\in\arg\max_{y\in\mathcal{S}(\bar{x})}F(\bar{x},y)$. Since $\mathcal{S}(\bar{x})$ is convex, and $F(\bar{x},y)$ is strongly concave in $y$, we must have $\bar{y}=y^{*}(\bar{x})$. Therefore, all accumulation points of $\{y_{k}^{*}(x_{k})\}$ equal $y^{*}(\bar{x})$. Then, the boundedness of $\{y_{k}^{*}(x_{k})\}$ implies that $\lim_{k\to\infty}y_{k}^{*}(x_{k})=y^{*}(\bar{x})$.
Finally, we show that
$\lim_{k\to\infty}z_{k}^{*}(x_{k})=y^{*}(\bar{x}).$
Since $z_{k}^{*}(x_{k})$ minimizes $\psi_{k}(x_{k},y_{k}^{*}(x_{k}),z)$ over $z\in Y$, we have
$\psi_{k}(x_{k},y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k}))\leq\psi_{k}(x_{k},y_{k}^{*}(x_{k}),y^{*}(x_{k})).$
Expanding this inequality gives
$\rho_{k}f(x_{k},z_{k}^{*}(x_{k}))+\frac{\sigma_{k}}{2}\|z_{k}^{*}(x_{k})\|^{2}-\sigma_{k}\langle y_{k}^{*}(x_{k}),z_{k}^{*}(x_{k})\rangle$
$\leq$ $\rho_{k}f(x_{k},y^{*}(x_{k}))+\frac{\sigma_{k}}{2}\|y^{*}(x_{k})\|^{2}-\sigma_{k}\langle y_{k}^{*}(x_{k}),y^{*}(x_{k})\rangle.$
Rearranging terms yields
$\rho_{k}\left(f(x_{k},z_{k}^{*}(x_{k}))-f(x_{k},y^{*}(x_{k}))\right)+\frac{\sigma_{k}}{2}\|z_{k}^{*}(x_{k})-y_{k}^{*}(x_{k})\|^{2}\leq\frac{\sigma_{k}}{2}\|y^{*}(x_{k})-y_{k}^{*}(x_{k})\|^{2}.$
Since $y^{*}(x_{k})\in\mathcal{S}(x_{k})$, it follows that $f(x_{k},z_{k}^{*}(x_{k}))-f(x_{k},y^{*}(x_{k})\geq 0$ and hence
$\|z_{k}^{*}(x_{k})-y_{k}^{*}(x_{k})\|^{2}\leq\|y^{*}(x_{k})-y_{k}^{*}(x_{k})\|^{2}.$
Because $\lim_{k\to\infty}y^{*}(x_{k})=y^{*}(\bar{x})=\lim_{k\to\infty}y_{k}^{*}(x_{k})$, we have
$\lim_{k\to\infty}\|y^{*}(x_{k})-y_{k}^{*}(x_{k})\|=0.$
Taking $k\rightarrow\infty$ in the above inequality yields
$\lim_{k\to\infty}\|z_{k}^{*}(x_{k})-y_{k}^{*}(x_{k})\|=0,$
and consequently,
$\lim_{k\to\infty}z_{k}^{*}(x_{k})=\lim_{k\to\infty}y_{k}^{*}(x_{k})=y^{*}(\bar{x}).$
This completes the proof.
$\square$